OpenStudy (anonymous):

Graph the line containing the given pair and find the slope (3,4),(-3,-2)

OpenStudy (anonymous):

first we have to find the slope m=-2-4/-3-3=-6/-6=1

OpenStudy (anonymous):

cool picture uploaded

OpenStudy (anonymous):

in order to graph this line we have to express this in slope intercept form y=mx + b we already found m and we can find b by substituting one of our coordinates substituting (3,4) we have 4=3+b b=1 thus our equation to graph is y=x+1 :D

OpenStudy (anonymous):

so,what is the slope again?

OpenStudy (anonymous):

which one is the slope?

OpenStudy (anonymous):

m=slope

OpenStudy (anonymous):

so 1?

OpenStudy (anonymous):

yep. let me know if there's anything else you need help with

OpenStudy (anonymous):

I need the slope as an integer or a fraction

OpenStudy (anonymous):

as an integer slope=1 as a fraction slope =1/1